If the point of the nail can be approximated as a circle with radius 2.00 x 10–3m, what is the pressure in magapascals extorted on the wall of a hammer strikes the the nail with a force of 104 N? Use 3.14 for pi
1
Expert's answer
2020-12-14T14:41:26-0500

Pressure ,"P= F\/A"

"A= \u03c0r\u00b2"

A = 3.14 ×(2×10-3m)= 0.00628m2

P=( 104/0.00628)=16560.5pa

P= 0.0165605Mpa